cos x°=sin20° also tan 20°­ = cot(x+30)° Sory but i am hopeless at this
Answer by tommyt3rd(5050)

for the first problem:
x=arccos(sin(20))=70
for the second:
cot(x)=1/tan(x)
so we can write
tan(x+30)=cot(20)
x+30=arctan(tan(x+30)=arctan(cot(20)
x+30=70
x=40